Diffusion Morning Report 10/30/03

ted berg tberg at snf.stanford.edu
Thu Oct 30 06:59:34 PST 2003


Good Morning All,
    Looking pretty good this morning.  Logs on Nitride and LTO are not 
updated.  Tystar is down for pull and clean.  Epi is still
glitching occasionally.

Diffusion 	Date: 	10/30/2003 	 
Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-Dirty Ox 	Up 	  	 
Tylan 8-SiGe 	Up 	  	12.6/30
Tylan 9-Poly 	Up 	  	6.35/30
Tylan 10-Nitride 	Up 	  	8.7/30
Tylan 11-BPSG-LTO 	Up 	  	27/30
Tylan 12-FGA 	Up 	  	 
Tystar 	down 	Pull and clean 	0/30
Doped Poly 	GONE 	  	 
TEOS 	yellow 	  	 
Tylan 16 	N/A 	  	 
Burn Box 	Up 	  	 
ZYGO 	Up 	  	 
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	Auto changeover working well 	Currently 480- B(c4) 
on line


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20031030/ebd5af9e/attachment.html>


More information about the diffusion mailing list